how do you set up the problem mix raisins cost that $5.00 per kg with peanuts that cost $3.80 a kg. How many kg of raisins should be mixed with 10kg of peanuts to obtain a mixture worth $4.00 per kg
Answer by checkley71(8403)

5x+3.8(10-x)=40 x being the kg amount of raisons.
5x+38-3.8x=40
5x-3.8x+38=40
1.2x=40-38
1.2x=2
x=2/1.2
x=1 2/3kg multiply this by 5=8 1/3 kg of raisons
10-x=8 1/3kg multiplied by 3.8=31 2/3 kg of peanuts
8 1/3+31 2/3= 40.